OpenStudy (maddytarin_98):

What are the solutions to the equation 3(x – 4)(x + 5) = 0? x = –4 or x = 5 x = 3, x = 4, or x = –5 x = 3, x = –4, or x = 5 x = 4 or x = –5

OpenStudy (rolldextro):

Set each term equal to zero (x-4)=0=) x=4 (x+5)=0 =) x=-5 The last choice "x = 4 or x = –5"